# Approaches to critical point theory via sequential and parametrized topological complexity

### v2

@@ 2. Sequential subspace complexities as index functions
Throughout this section we choose and fix some $r\in\mathbb{N}$ with $r\geq 2$.
@@
The following lemma is an immediate consequence of the assertions of [Rud10, Remark 3.2] and an analogue of [Far06, Lemma 18.1]. For the reader’s convenience, we provide an explicit proof.
Lemma 2.2.
@@
    $A$ admits a continuous sequential motion planner.
@@
$h\colon A\times[0,1]\to X^{r},$
$h(a,t)=\Big{(}s(a)(0),s(a)\big{(}\tfrac{1-t}{r-1}\big{)},s(a)\big{(}\tfrac{2(1-t)}{r-1}\big{)},\dots,s(a)\big{(}\tfrac{(r-2)(1-t)}{r-1}\big{)},s(a)(1-t)\Big{)},$
@@
(2.1) $(s(a))(t)=\begin{cases}h_{j-1}(a,2((r-1)t-j+1))&\text{if }t\in\Big{[}\frac{j-1}{r-1},\frac{j-\frac{1}{2}}{r-1}\Big{)},\\ h_{j}(a,2(j-(r-1)t))&\text{if }t\in\Big{[}\frac{j-\frac{1}{2}}{r-1},\frac{j}{r-1}\Big{]},\end{cases}\qquad\forall j\in\{1,2,\dots,r-1\}.$
@@
To establish the continuity property of $\nu(A)=\operatorname{\mathsf{TC}}_{r,X}(A)$, we need to make an additional assumption on the space under consideration. To clarify the type of spaces that we will consider, we recall some terminology.
Definition 2.3.
A topological space $Y$ is called an absolute neighborhood retract (ANR) if it is metrizable and satisfies the following condition:
Let $X$ be a metrizable space and let $A\subset X$ be closed. Then for any continuous map $f\colon A\to Y$ there are an open neighborhood $U\subset X$ of $A$ and a continuous map $F\colon U\to Y$ with $F|_{A}=f$.
Remark 2.4.
Note that this definition coincides with the one used in [CLOT03] and [Pal66a], but differs from the one in [Hu65], which we will occasionally use as a reference. More precisely, our notion of an ANR coincides with the definition of an absolute neighborhood extensor (ANE) in [Hu65].
@@
One checks without difficulties that $H$ is well-defined and continuous. Since $X$ is an ANR, $X^{r}$ is an ANR as well. Hence, there exist an open neighborhood $V\subset X^{r}\times[0,1]$ of $Q$ and a continuous map $K\colon V\to X^{r}$ with $K|_{Q}=H$. Since $[0,1]$ is compact, there exists an open subset $U\subset X^{r}$ with $U\times[0,1]\subset V$ and $A\subset U$. By definition of the map, $K|_{U\times[0,1]}$ is a deformation of $U$ with $K(x,1)=d(x)$ for each $x\in U$. Since $X$ is by assumption metrizable, it is a Hausdorff space, so it follows from Lemma 2.2.a) that $U$ admits a continuous sequential motion planner. ∎
Using the previous lemma, we can now establish that sequential subspace complexities indeed form index functions.
@@
    For all $A\subset B\subset X^{r}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(A)\leq\operatorname{\mathsf{TC}}_{r,X}(B).$
@@
    For all $A,B\subset X^{r}$ it holds that
    $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,X}(A)+\operatorname{\mathsf{TC}}_{r,X}(B).$
@@
  1. a)
    This is obvious.
  2. b)
    This follows immediately from the definition of $\operatorname{\mathsf{TC}}_{r,X}$.
@@
    Let $k:=\operatorname{\mathsf{TC}}_{r,X}(A)$. Then by definition of $\operatorname{\mathsf{TC}}_{r,X}(A)$ there are $k$ open domains $U_{1},\dots,U_{k}\subset X^{r}$ of continuous sequential motion planners, such that $A\subset\bigcup_{i=1}^{k}U_{i}$. Put $U:=\bigcup_{i=1}^{k}U_{i}$. Since $A\subset U$, we derive from a) that $k\leq\operatorname{\mathsf{TC}}_{r,X}(U)$. However, since $U_{1},\dots,U_{k}$ form an open cover of $U$ for which each element admits a continuous sequential motion planner, it follows that $\operatorname{\mathsf{TC}}_{r,X}(U)\leq k$. Together, the two inequalities show the claim.
@@
    It follows immediately from a) that $\max\{\operatorname{\mathsf{TC}}_{r,X}(A),\operatorname{\mathsf{TC}}_{r,X}(B)\}\leq\operatorname{\mathsf{TC}}_{r,X}(A\cup B)$, so it only remains to show the opposite inequality. Let $m:=\operatorname{\mathsf{TC}}_{r,X}(A)$ and $n:=\operatorname{\mathsf{TC}}_{r,X}(B)$ and assume w.l.o.g. that $m\geq n$. Let $U_{1},\dots,U_{m},V_{1},\dots,V_{n}\subset X^{r}$ be open subsets, such that there are continuous sequential motion planners $s_{i}\colon U_{i}\to PX$ and $t_{j}\colon V_{j}\to PX$ for all $i\in\{1,2,\dots,m\}$ and $j\in\{1,2,\dots,n\}$ and such that
    $A\subset\bigcup_{i=1}^{m}U_{i},\qquad B\subset\bigcup_{i=1}^{n}V_{j}.$
    If $m>n$, then put $V_{j}:=\emptyset$ for all $j\in\{m+1,\dots,n\}$. Since $X$ is normal, $X^{r}$ is normal as well, so there are open subsets $U_{0},V_{0}\subset X^{r}$ with $A\subset U_{0}$, $B\subset V_{0}$ and $U_{0}\cap V_{0}=\emptyset$. Put $U_{i}^{\prime}:=U_{0}\cap U_{i}$ and $V_{i}^{\prime}:=V_{0}\cap V_{i}$ for all $i\in\{1,2,\dots,m\}$. Then for each $i$ we obtain a continuous sequential motion planner
    $U^{\prime}_{i}\cup V_{i}^{\prime}\to PX,\qquad x\mapsto\begin{cases}s_{i}(x)&\text{if }x\in U_{i}^{\prime},\\ t_{i}(x)&\text{if }x\in V_{i}^{\prime}.\end{cases}$
    Since $A\cup B\subset\bigcup_{i=1}^{m}(U_{i}^{\prime}\cup V_{i}^{\prime})$, this shows that $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)\leq m=\max\{\operatorname{\mathsf{TC}}_{r,X}(A),\operatorname{\mathsf{TC}}_{r,X}(B)\}$.
@@
Put $k:=\operatorname{\mathsf{TC}}_{r,X}(\Phi_{1}(A))$ and let $U_{1},\dots,U_{k}\subset X^{r}$ be open subsets with $\Phi_{1}(A)\subset\bigcup_{j=1}^{k}U_{j}$, such that each $U_{j}$ admits a continuous sequential motion planner $s_{j}\colon U_{j}\to PX$. Since $X$ is an ANR, $X$ is metrizable, hence normal. As a closed subspace of $X^{r}$, $\Phi_{1}(A)$ is normal as well. Thus, see e.g. [CLOT03, Theorem A.1.(6)], we can find closed subsets $B_{j}\subset U_{j}$ for $j\in\{1,2,\dots,k\}$ with $\Phi_{1}(A)=\bigcup_{j=1}^{k}B_{j}$. Put $B_{j}^{\prime}:=\Phi_{1}^{-1}(B_{j})$ for each $j\in\{1,2,\dots,k\}$. Then $A\subset\Phi_{1}^{-1}(\Phi_{1}(A))\subset\bigcup_{j=1}^{k}B_{j}^{\prime}$. Moreover, for each $j$ we obtain a continuous motion planner $s_{j}^{\prime}\colon B_{j}^{\prime}\to PX$ as follows: Denote the components of $\Phi$ by $\phi_{j}\colon A\times[0,1]\to X$, $j\in\{1,2,\dots,r\}$, such that $\Phi(x,t)=(\phi_{1}(x,t),\dots,\phi_{r}(x,t))$. Given $x=(x_{0},x_{1},\dots,x_{r-1})\in B_{j}^{\prime}$, $j\in\{1,2,\dots,r-1\}$ and $t\in\left[\frac{j-1}{r-1},\frac{j}{r-1}\right]$, we put
(2.3) $(s_{j}^{\prime}(x))(t):=\begin{cases}\phi_{j-1}(x,3(r-1)t-3j+3)&\text{if }t\in\Big{[}\frac{j-1}{r-1},\frac{j-\frac{2}{3}}{r-1}\Big{)},\\ (s_{j}(\Phi_{1}(x)))\left(3t+\frac{1-2j}{r-1}\right)&\text{if }t\in\Big{[}\frac{j-\frac{2}{3}}{r-1},\frac{j-\frac{1}{3}}{r-1}\Big{)},\\ \phi_{j}(x,3j-3(r-1)t)&\text{if }t\in\Big{[}\frac{j-\frac{1}{3}}{r-1},\frac{j}{r-1}\Big{]}.\end{cases}$
@@
Definition 2.9.
  1. a)
    Let $X$ be a topological space, $F\colon X\to\mathbb{R}$ and $\varphi\colon X\to X$. For each $\mu\in\mathbb{R}$ we put
    $\mathcal{C}_{\mu}(\varphi,F):=\pi_{0}(F_{\mu}\cap\mathrm{Fix\;}\varphi),$
    i.e. the set of connected components of $F_{\mu}\cap\mathrm{Fix\;}\varphi$.
  2. b)
    Let $M$ be a smooth Banach manifold and $F\in C^{1}(M)$. For each $\mu\in\mathbb{R}$ we put
    $\mathcal{C}_{\mu}(F):=\pi_{0}(F_{\mu}\cap\operatorname{Crit}F).$
By Corollary 2.8 we can apply Theorem 1.4 to the index function $\operatorname{\mathsf{TC}}_{r,X}(\cdot)$, to obtain the following result.
@@
    This is an immediate consequence of combining Corollary 2.8 and Theorem 1.4.
@@
By [Pal66b, Theorem 4.4], see also [Str08, Lemma II.3.9], the function $F$ admits a pseudo-gradient $V\colon X^{r}\to TX^{r}$ in the sense of Definition 1.5. Let $\phi_{1}\colon X^{r}\to X^{r}$ be the time-1 map of the flow of $-V$, which is well-defined by the completeness of $(X^{r},g^{r})$. The flow yields a homotopy from $\phi_{1}$ to the identity map. By Theorem 1.10, $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D). Moreover, by Lemma 1.9, $\mathcal{C}_{\mu}(F)$ is finite for all $\mu\in\mathbb{R}$. Thus, the claim follows directly from Theorem 2.10.b), since by definition of $\phi_{1}$ and by Remark 1.6.(3) it holds that $\mathrm{Fix\;}\phi_{1}=\mathrm{sing}\,V=\operatorname{Crit}F$. ∎
@@ 3. Navigation functions and sequential subspace complexities
    By definition of an $r$-navigation function $F\colon X^{r}\to\mathbb{R}$, its global minimum is 00 and it holds that
@@
[…]
@@
[…]
(3.2) $\operatorname{\mathsf{TC}}(M)=\operatorname{\mathsf{TC}}(F^{\lambda})\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1.$
@@
$\alpha(M)=\#\{(x,y)\in\operatorname{Crit}F\mid x\neq y\}=\#\{(x,y)\in\operatorname{Crit}F\mid F(x,y)>0\}.$
Since $\alpha(M)$ is by assumption finite, $F_{\mu}\cap\operatorname{Crit}F$ is finite for each $\mu>0$. Thus, each $C\in\mathcal{C}_{\mu}(F)$ consists of a single point, which immediately yields $\operatorname{\mathsf{TC}}_{M}(C)=1$. We derive from (3.2) that
$\operatorname{\mathsf{TC}}(M)$$\leq\#\{\mu\in(0,\lambda]\mid F_{\mu}\cap\operatorname{Crit}F\neq\emptyset\}+1$
#$\leq\#\{(x,y)\in\operatorname{Crit}F\mid F(x,y)>0\}+1=\alpha(M)+1.$
The claim immediately follows. ∎
So far, our guideline was to estimate the number of critial points of functions from below in terms of sequential subspace complexities. In the following, we want to go the opposite way and show in an example how to use $r$-navigation functions to estimates values of sequential TCs from above.
@@
They obtained this value as a lower bound for $\operatorname{\mathsf{TC}}_{r}(M)$ by a standard cohomological computation and as an upper bound using an inequality for sequential TCs of product spaces in terms of the sequential TCs of its factors. Here, we want to show that the upper bound can also be obtained by using $r$-navigation functions.
For $n\in\mathbb{N}$ we consider the $n$-sphere as
$\mathbb{S}^{n}=\{x\in\mathbb{R}^{n+1}\mid\|x\|=1\},$
[…]
@@
Let $x=(x_{1},\dots,x_{r})\in\operatorname{Crit}f_{n}$. We will prove by induction over $i$ that $\{x_{1},x_{i}\}$ is linearly dependent for each $i\in\{2,3,\dots,r\}$. Using (3.1), we first compute for each $v_{1}\in T_{x_{1}}\mathbb{S}^{n}$ that
$0=DF_{x}(v_{1},0,\dots,0)=2\left<x_{1}-x_{2},v_{1}\right>=2\left<x_{2},v_{1}\right>.$
Thus, $T_{x_{1}}\mathbb{S}^{n}\subset\{x_{2}\}^{\perp}$ and since $T_{x_{1}}\mathbb{S}^{n}=\{x_{1}\}^{\perp}$ it follows for dimensional reasons that $\{x_{1}\}^{\perp}=\{x_{2}\}^{\perp}$, so $x_{1}$ and $x_{2}$ are linearly dependent.
Assume as induction hypothesis that $x_{1}$ and $x_{j}$ are linearly dependent for all $j\in\{2,3,\dots,i\}$ for some $i\geq 2$. We compute for each $v_{i}\in T_{x_{i}}\mathbb{S}^{n}$ that
[…]
since $x_{i},v_{i}\right>=0$. In analogy with the case of $i=2$, this shows that $T_{x_{i}}\mathbb{S}^{n}\subset\{x_{i-1}+x_{i+1}\}^{\perp}$ and consequently that $x_{i}$ and $x_{i-1}+x_{i+1}$ are linearly dependent, i.e. there are $\lambda_{1},\lambda_{2}\in\mathbb{R}$ with $(\lambda_{1},\lambda_{2})\neq(0,0)$ and
$\lambda_{1}x_{i}+\lambda_{2}(x_{i-1}+x_{i+1})=0.$
If $\lambda_{2}=0$, then it would immediately follow that $x_{i}=0$, which is a contradiction. Hence, $\lambda_{2}\neq 0$ and we can rewrite the equation as $x_{i+1}=-\frac{\lambda_{1}}{\lambda_{2}}x_{i}-x_{i-1}$. By our induction hypothesis, both $x_{i}$ and $x_{i-1}$ are linearly dependent of $x_{1}$, so it follows immediately that $x_{i+1}$ is linearly dependent of $x_{1}$ as well, which completes our induction. Hence, we have shown that $x_{i}\in\mathbb{R}x_{1}$ for each $i$ and since $x_{1},\dots,x_{r}\in\mathbb{S}^{n}$, this shows that $x_{i}\in\{x_{1},-x_{1}\}$ for each $i\in\{1,2,\dots,r\}$. Conversely, a straightforward calculation which we omit shows that any such point is indeed a critical point of $f_{n}$. ∎
@@
$f_{n}(x_{1},\dots,x_{r})=4m,\quad\text{where }\ \ m=\#\{i\in\{2,3,\dots,r\}\mid x_{i}\neq x_{1}\}.$
Thus, the set of critical values of $f_{n}$ is given as
$f_{n}(\operatorname{Crit}f_{n})=\{0,4,8,\dots,4(r-1)\}.$
For $n_{1},n_{2},\dots,n_{k}\in\mathbb{N}$ we let again $M=\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\mathbb{S}^{n_{k}}$ and consider the function
$F\colon M^{r}\to\mathbb{R},\quad F((x_{11},\ldots,x_{1k}),\ldots,(x_{r1},\ldots,x_{rk}))=\sum_{j=1}^{k}f_{n_{j}}(x_{1j},x_{2j},\dots,x_{rj}).$
@@
$\operatorname{Crit}F$ =$=\left\{\left((x_{11},\dots,x_{1k}),\dots,(x_{r1},\dots,x_{rk})\right)\in M^{r}\ \middle|\ (x_{1j},\dots,x_{rj})\in\operatorname{Crit}f_{n_{j}}\ \forall j\in\{1,2,\dots,k\}\right\}$
=$=\left\{\left((x_{11},\dots,x_{1k}),\dots,(x_{r1},\dots,x_{rk})\right)\in M^{r}\ \middle|\ x_{ij}\in\{x_{1j},-x_{1j}\}\ \forall i\in\{1,\dots,r\},\ j\in\{1,\dots,k\}\right\},$

### v4

@@ 2. Sequential subspace complexities as index functions
Throughout this section we fix some $r\in\mathbb{N}$ with $r\geq 2$.
@@
The $r$-th sequential topological complexity of $X$, denoted by $\operatorname{\mathsf{TC}}_{r}(X)$, is the sectional category of $p_{r}$ and was introduced by Rudyak in [Rud10] under the name of higher topological complexity. In the following definition, we introduce a relative version of this notion, which was first investigated by Daundkar in [Dau24].
@@
Remark 2.2.
Let $X$ be a path-connected topological space.
  1. (1)
    Our definition of $\operatorname{\mathsf{TC}}_{r,X}(A)$ slightly differs from the one in [Dau24] as Daundkar considers covers of $A$ by subsets of $A$ which are open in $A$ with respect to the subspace topology. By standard arguments, one shows that both notions coincide if $X$ is an absolute neighborhood retract and if $A$ is closed in $X^{r}$.
  2. (2)
    In terms of Definition 2.1 the $r$-th sequential topological complexity of $X$ is given as $\operatorname{\mathsf{TC}}_{r}(X)=\operatorname{\mathsf{TC}}_{r,X}(X^{r})$. However, we note that if $B\subseteq X$ is a subspace then the $r$-th sequential subspace complexity $\operatorname{\mathsf{TC}}_{r,X}(B^{r})$ can be different from the intrinsic $r$-th sequential topological complexity $\operatorname{\mathsf{TC}}_{r}(B)$. This can already be seen in the case $r=2$ taking $X=\mathbb{S}^{2}$ and $B=\mathbb{S}^{1}$ where $\mathbb{S}^{1}$ is embedded as the equator. It is well-known that $\operatorname{\mathsf{TC}}_{2}(\mathbb{S}^{1})=2$, but since $\mathbb{S}^{1}$ is contractible in $\mathbb{S}^{2}$ we obtain $\operatorname{\mathsf{TC}}_{2,\mathbb{S}^{2}}(\mathbb{S}^{1}\times\mathbb{S}^{1})=1$.
  3. (3)
    Sequential topological complexity is motivated by an idea from robot motion planning. Consider a robot working in the configuration space $X$ equipped with a motion planning algorithm, so that for an arbitrary $r$-tuple of points, the robot is able to find a path visiting all $r$ points of the tuple in the given order. $\operatorname{\mathsf{TC}}_{r}(X)$ describes the minimal number of discontinuities of such an algorithm. If one only wants the robot to find its way for a certain subset of point tuples $A\subset X^{r}$, but allows it to move inside all of $X$, the number $\operatorname{\mathsf{TC}}_{r,X}(A)$ expresses the minimal number of discontinuities of such an algorithm.
@@
The following lemma is an immediate consequence of the assertions of [Rud10, Remark 3.2] and an analogue of [Far06, Lemma 18.1], which has originally been shown in [Dau24]. For the reader’s convenience, we include its statement and proof.
Lemma 2.3 ([Dau24, Lemma 2.4]).
@@
    AA admits a continuous sequential motion planner.
@@
$h\colon A\times[0,1]\to X^{r},\quad h(a,t)=\Big(s(a)(0),s(a)\big(\tfrac{1-t}{r-1}\big),,\dots,s(a)\big(\tfrac{(r-2)(1-t)}{r-1}\big),s(a)(1-t)\Big),$
@@
(2.1) $s(a)(t)=\begin{cases}h_{j-1}(a,2((r-1)t-j+1))&\text{if }t\in\Big[\frac{j-1}{r-1},\frac{j-\frac{1}{2}}{r-1}\Big),\\ h_{j}(a,2(j-(r-1)t))&\text{if }t\in\Big[\frac{j-\frac{1}{2}}{r-1},\frac{j}{r-1}\Big],\end{cases}\qquad\forall j\in\{1,2,\dots,r-1\}.$
@@
To establish the continuity property of $\nu(A)=\operatorname{\mathsf{TC}}_{r,X}(A)$, we need to make an additional assumption on the space under consideration. In the following results, we will assume that the considered space is an absolute neighborhood retract (ANR), see e.g. [CLOT03, p. 287] or [Pal66a] for a definition.
@@
One checks without difficulties that $H$ is well-defined and continuous. Since $X$ is an ANR, $X^{r}$ is an ANR as well. Hence, there exist an open neighborhood $V\subset X^{r}\times[0,1]$ of $Q$ and a continuous map $K\colon V\to X^{r}$ with $K|_{Q}=H$. Since $[0,1]$ is compact, there exists an open subset $U\subset X^{r}$ with $U\times[0,1]\subset V$ and $A\subset U$. By definition of the map, $K|_{U\times[0,1]}$ is a deformation of $U$ with $K(x,1)=d(x)$ for each $x\in U$. Since $X$ is by assumption metrizable, it is a Hausdorff space, so it follows from Lemma 2.3 that $U$ admits a continuous sequential motion planner. ∎
Using the previous lemma, we can now assert that sequential subspace complexities form index functions.
@@
    For all $A\subset B\subset X^{r}$, it holds that $\operatorname{\mathsf{TC}}_{r,X}(A)\leq\operatorname{\mathsf{TC}}_{r,X}(B)$.
@@
    For all $A,B\subset X^{r}$ it holds that $\operatorname{\mathsf{TC}}_{r,X}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,X}(A)+\operatorname{\mathsf{TC}}_{r,X}(B)$.
@@
  1. a) and b)
    These statements follow immediately from the definition of $\operatorname{\mathsf{TC}}_{r,X}$.
@@
    Let $k:=\operatorname{\mathsf{TC}}_{r,X}(A)$. Then by definition of $\operatorname{\mathsf{TC}}_{r,X}(A)$ there are $k$ open domains $U_{1},\dots,U_{k}\subset X^{r}$ of continuous sequential motion planners, such that $A\subset\bigcup_{i=1}^{k}U_{i}$. Put $U:=\bigcup_{i=1}^{k}U_{i}$. Since $A\subset U$, we derive from a) that $k\leq\operatorname{\mathsf{TC}}_{r,X}(U)$. However, since $U_{1},\dots,U_{k}$ form an open cover of $U$ for which each element admits a continuous sequential motion planner, it follows that $\operatorname{\mathsf{TC}}_{r,X}(U)\leq k$.
@@
    This follows by standard arguments, see e.g. [Fox41, p.336] for the corresponding result for Lusternik-Schnirelmann category.
@@
Put $k:=\operatorname{\mathsf{TC}}_{r,X}(\Phi_{1}(A))$ and let $U_{1},\dots,U_{k}\subset X^{r}$ be open subsets with $\Phi_{1}(A)\subset\bigcup_{j=1}^{k}U_{j}$, such that each $U_{j}$ admits a continuous sequential motion planner $s_{j}\colon U_{j}\to PX$. Since $X$ is an ANR, $X$ is metrizable, hence normal. As a closed subspace of $X^{r}$, $\Phi_{1}(A)$ is normal as well. Thus, see e.g. [CLOT03, Theorem A.1.(6)], we can find closed subsets $B_{j}\subset U_{j}$ for $j\in\{1,2,\dots,k\}$ with $\Phi_{1}(A)=\bigcup_{j=1}^{k}B_{j}$. Put $B_{j}^{\prime}:=\Phi_{1}^{-1}(B_{j})$ for each $j\in\{1,2,\dots,k\}$. Then $A\subset\Phi_{1}^{-1}(\Phi_{1}(A))\subset\bigcup_{j=1}^{k}B_{j}^{\prime}$. Moreover, for each $j$ we obtain a continuous motion planner $s_{j}^{\prime}\colon B_{j}^{\prime}\to PX$ as follows: Denote the components of $\Phi$ by $\phi_{j}\colon A\times[0,1]\to X$, $j\in\{1,2,\dots,r\}$, such that $\Phi(x,t)=(\phi_{1}(x,t),\dots,\phi_{r}(x,t))$. Given $x=(x_{0},x_{1},\dots,x_{r-1})\in B_{j}^{\prime}$, $\ell\in\{1,2,\dots,r-1\}$ and $t\in\left[\frac{\ell-1}{r-1},\frac{\ell}{r-1}\right]$, we put
[…]
@@
Let $X$ be a topological space, $F\colon X\to\mathbb{R}$ and $\varphi\colon X\to X$. For each $\mu\in\mathbb{R}$ we let $\mathcal{C}_{\mu}(\varphi,F)$ denote the set of connected components of $F_{\mu}\cap\mathrm{Fix\;}\varphi$. In the case that $X$ is a smooth Banach manifold and that $F$ is differentiable, let $\mathcal{C}_{\mu}(F)$ denote the set of connected components of $F_{\mu}\cap\operatorname{Crit}F$.
The following result is a generalization of [Far08, Theorem 4.32].
@@
    This is an immediate consequence of Corollary 2.7 and Theorem 1.4.
@@
By [Pal66b, Theorem 4.4], see also [Str08, Lemma II.3.9], the function $F$ admits a pseudo-gradient $V\colon X^{r}\to TX^{r}$ in the sense of Definition 1.5. Let $\phi_{1}\colon X^{r}\to X^{r}$ be the time-1 map of the flow of V-V, which is well-defined by the completeness of $(X^{r},g^{r})$. The flow yields a homotopy from $\phi_{1}$ to the identity map. By Theorem 1.10, $F$ is a Lyapunov function for $\phi_{1}$ and $(\phi_{1},F)$ satisfies condition (D). Moreover, by Lemma 1.9, $\mathcal{C}_{\mu}(F)$ is finite for all $\mu\in\mathbb{R}$. Thus, the claim follows directly from Theorem 2.8.b), since by definition of $\phi_{1}$ and by Remark 1.6.(3) it holds that $\mathrm{Fix\;}\phi_{1}=\mathrm{sing}\,V=\operatorname{Crit}F$. ∎
@@ 3. Navigation functions and sequential subspace complexities
    By definition of an $r$-navigation function $F\colon X^{r}\to\mathbb{R}$, its global minimum is 0 and it holds that
@@
  3. (3)
    The motivation for the introduction of navigation functions is the wish to carry out motion planning along paths following gradient flow lines of a differentiable function. This idea was first established by D. Koditschek and E. Rimon in [KR90] who considered motion planning to a fixed end point along gradient flow lines. The idea behind a $2$-navigation function is to use an analogous method for flexible start and end points. This approach is explained in greater detail in [Mes24, Section 4.1].
@@
Let $n,r\in\mathbb{N}$ with $r\geq 2$, let $M$ be a compact submanifold of $\mathbb{R}^{n}$ (without boundary) and let $\|\cdot\|$ denote the Euclidean norm on $\mathbb{R}^{n}$. Then
@@
If $\alpha(M)$ is infinite, the statement is evident, so we assume throughout the following that $\alpha(M)$ is finite. Consider the function $F\colon M\times M\to\mathbb{R}$, $F(x,y)=\|x-y\|^{2}$, where $\|\cdot\|$ denotes the Euclidean metric on $\mathbb{R}^{n}$. Since $M$ is compact, $F$ attains its maximum $\lambda$. As explained in Example 3.4, $F$ is a $2$-navigation function, so it follows from Theorem 3.3 in the case of $r=2$ that
(3.2) $\operatorname{\mathsf{TC}}(M)=\operatorname{\mathsf{TC}}(F^{\lambda})\leq\sum_{\mu\in(0,\lambda]}\max\{\operatorname{\mathsf{TC}}_{2,M}(C)\mid C\in\mathcal{C}_{\mu}(F)\}+1.$
@@
(3.3) $\alpha(M)=\#\{\{x,y\}\subset M\mid(x,y)\in\operatorname{Crit}F,\ F(x,y)>0\}.$
Since $\alpha(M)$ is by assumption finite, $F_{\mu}\cap\operatorname{Crit}F$ is finite for each $\mu>0$. Thus, each $C\in\mathcal{C}_{\mu}(F)$ consists of a single point, which immediately yields $\operatorname{\mathsf{TC}}_{2,M}(C)=1$. Thus, we derive from (3.2) that
(3.4) $\operatorname{\mathsf{TC}}(M)\leq\#\{\mu\in(0,\lambda]\mid F_{\mu}\cap\operatorname{Crit}F\neq\emptyset\}+1.$
While for any $(x,y)\in\operatorname{Crit}F$ we also have $(y,x)\in\operatorname{Crit}F$, both of these points lie in the same level set of $F$. Thus, it follows from (3.3) and (3.4) that $\operatorname{\mathsf{TC}}(M)\leq\alpha(M)+1$. ∎
So far, our guideline was to estimate the number of critical points of functions from below in terms of sequential subspace complexities. In the following, we want to go the opposite way and show in an example how to use $r$-navigation functions to estimates values of sequential TCs from above.
@@
They obtained this value as a lower bound of $\operatorname{\mathsf{TC}}_{r}(M)$ by a standard cohomological computation and as an upper bound using an inequality for sequential TCs of product spaces in terms of the sequential TCs of its factors. Here, we want to show that the upper bound can also be obtained by using $r$-navigation functions.
For $n\in\mathbb{N}$ we consider the $n$-sphere as $\mathbb{S}^{n}=\{x\in\mathbb{R}^{n+1}\mid\|x\|=1\}$, so that for $x\in\mathbb{S}^{n}$ we have $T_{x}\mathbb{S}^{n}=\{v\in\mathbb{R}^{n+1}\mid\left<x,v\right>=0\}=\{x\}^{\perp},$ where $\left$$\cdot$$\cdot\right$> and $\|\cdot\|$ again denote the Euclidean inner product and the Euclidean norm, respectively. We further consider the $r$-navigation function
@@
Let $x=(x_{1},\dots,x_{r})\in\operatorname{Crit}f_{n}$. Using (3.1), we first compute for each $v_{1}\in T_{x_{1}}\mathbb{S}^{n}$ that
$0=D(f_{n})_{x}(v_{1},0,\dots,0)=2\left<x_{1}-x_{2},v_{1}\right>=-2\left<x_{2},v_{1}\right$>.
Thus, $T_{x_{1}}\mathbb{S}^{n}\subset\{x_{2}\}^{\perp}$ and since $T_{x_{1}}\mathbb{S}^{n}=\{x_{1}\}^{\perp}$ it follows for dimensional reasons that $\{x_{1}\}^{\perp}=\{x_{2}\}^{\perp}$, so $x_{1}$ and $x_{2}$ are linearly dependent. A straightforward inductive argument shows that the set $\{x_{1},x_{i}\}$ is linearly dependent for $i\in\{2,3\dots,r\}$. Hence, $x_{i}\in\{x_{1},-x_{1}\}$ for each $i\in\{1,2,\dots,r\}$. Conversely, a direct calculation which we omit shows that any such point is indeed a critical point of $f_{n}$. ∎
@@
$f_{n}(x_{1},\dots,x_{r})=4m,\quad\text{where }\ \ m=\#\{i\in\{2,3,\dots,r\}\mid x_{i}\neq x_{i-1}\}.$
The set of critical values of $f_{n}$ is thus given as $f_{n}(\operatorname{Crit}f_{n})=\{0,4,\dots,4(r-1)\}.$ For integers $n_{1},\dots,n_{k}\in\mathbb{N}$ we let again $M=\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\mathbb{S}^{n_{k}}$ and consider the function
$F\colon M^{r}\to\mathbb{R},\quad F((x_{11},\ldots,x_{1k}),\ldots,(x_{r1},\ldots,x_{rk}))=\sum_{j=1}^{k}f_{n_{j}}(x_{1j},\dots,x_{rj}).$
@@
$\operatorname{Crit}F=\left\{\left((x_{11},\dots,x_{1k}),\dots,(x_{r1},\dots,x_{rk})\right)\in M^{r}\ \middle|\ x_{ij}\in\{x_{1j},-x_{1j}\}\ \forall i\in\{2,\dots,r\},\ j\in\{1,\dots,k\}\right\},$